A swimming pool filled with water has dimensions of 5.01 m × 10.2 m × 1.76 m. water has density ρ 1.00 103 kg/m) and sp c heat C-4186 C HINT (a) Find the mass (in kg) of water in the pool. kg (b) Find the thermal energy (in 3) required to heat the pool water from 15,8°C to 26.6c. t the pool water from 15.8°C to 26.6°c (c) Calculate the cost (in dollars) of heating the pool from 15.8°c to 26.6°C if electrical energy costs $0.115 per kalowatt-hour
Answer #1

a) The mass of water is = volume times density

M=5.01\times 10.2\times 1.76 \times10^3=89939.5\, kg

b) Energy required to heat the water is

E=M c \Delta T=89939.5\times4186\times (26.6 - 15.8)=4.066\times 10^9\,J

c) Energy is KiloWatt Hour is

E=4.066\times 10^9/(3.6\times 10^6)=1129.46\, kWh

Cost required to heat the water is

cost=E\times 0.115=\$ 129.89

cost is 129.89 dollars